OpenStudy (anonymous):

find the limit as x approaches 0 of 4x cosx/sin x

OpenStudy (anonymous):

\[4\lim_{x \rightarrow 0} \frac{ x }{ \sin x } * \lim_{x \rightarrow0} \cos x\]

OpenStudy (anonymous):

does that help?

OpenStudy (anonymous):

Kind of but then you plub in 0 for x and cos x cancels out?

OpenStudy (anonymous):

remember learning about the limit of (sin(x))/x as x->0

OpenStudy (anonymous):

no i didnt learn much about trig functions